# PTR: A Pre-trained Language Model for Trajectory Recovery

### v1

@@ 5.6. Scalibility Analysis (RQ2)
PTR 81.76 84.31 88.38 187.4 366.2
Porto w/o explicit trajectory prompt 66.13 66.19 81.98 157.9 314.6
w/o implicit trajectory prompt 66.18 66.22 82.01 149.6 298.1
w/o trajectory prompt 65.52 65.89 81.51 171.0 339.4
w/o road network 66.11 66.27 81.90 150.7 302.8
w/o reference tokens 65.89 65.07 80.57 165.4 327.7
PTR 66.40 66.52 82.14 141.9 294.6
5.7. Ablation Study (RQ3)
To evaluate the effects of various components in PTR, we conduct ablation experiments on the Chengdu and Porto datasets with sampling intervals of 2 minutes. We compare PTR with the following five variants.
  • w/o explicit trajectory prompt: We remove the explicit trajectory prompt.
  • w/o implicit trajectory prompt: We use the ‘[MASK]’ token in the BERT to represent the missing location.
  • w/o trajectory prompt: We remove both the explicit and implicit trajectory prompt, and the missing points are represented by the ‘[MASK]’ token in BERT.
As shown in Table 3, each component of PTR is essential. The explicit trajectory prompt provides task and trajectory-related textual information, enabling PLMs to effectively understand the trajectory recovery tasks. Omitting this component leads to a performance decrease. When the implicit trajectory prompt is removed, many missing locations are replaced by the ‘[MASK]’ token in BERT, limiting the model’s ability to capture external factors for missing points, which results in suboptimal performance.
For the trajectory point embedder, excluding road network for observed points causes the model to rely solely on latitude and longitude, neglecting the relationship between trajectories and the road network, which demonstrates that road network is crucial for accurately modeling trajectory information. Additionally, removing reference tokens results in a performance decline, as PLMs are unable to directly interpret trajectory data.
@@ 6. Conclusion
In this paper, we propose PTR, a novel framework for trajectory recovery leveraging PLM. PTR efficiently handles trajectory recovery even with limited dense trajectory data, demonstrating strong scalability. The model is trained on datasets with various sparse trajectories, enabling it to effectively generalize to different sampling intervals. To model external factors, we introduce an implicit trajectory prompt to model road conditions, which provides rich information for missing trajectory points. We also develop a trajectory embedder to capture the spatial-temporal dependencies of sparse trajectories and convert them into a format interpretable by PLMs. Experimental results on two datasets with three different sampling intervals validate the efficacy and scalability of the proposed model.
Appendix A Dataset
Table 4. Dataset Description.
Types Chengdu Porto
Sampling interval 15s 15s
# Trajectory 118,354 322,079
# Road segment 2504 2224
Latitude range 30.655 $\sim$ 30.727 41.142 $\sim$ 41.174
Longitude range 104.043 $\sim$ 104.129 -8.652 $\sim$ -8.578
Length / Width 8.22 km / 8.00 km 6.19 km / 3.56 km
Appendix B Baseline Setting
We choose the following ten methods as baselines, including five free space trajectory recovery and five map-matched trajectory recovery.
B.1. Free-space Trajectory Recovery
@@
B.2. Map-matched Trajectory Recovery
@@
  • T2vec (Li et al., 2018) + Decoder is a deep learning model for trajectory similarity learning. We use its encoder to embed the sparse trajectory and use the decoder of the MTrajRec to recover the trajectory in an autoregressive manner.
  • T3s (Yang et al., 2021) + Decoder uses LSTM and attention mechanisms to encode sparse trajectory data effectively. Then, the MTrajRec decoder is utilized to recover the trajectory.
@@
Appendix C Evaluation Metrics
@@
Appendix D Appendix Experiments
D.1. Covergence Rate
Figure 5. Convergence rate of PTR and fine-tuning for three sampling intervals on Chengdu.
Figure 6. Convergence rate of PTR and fine-tuning for three sampling intervals on Porto.
We evaluate the convergence speed of our model on the Porto and Chengdu datasets. As illustrated in Figure 5 and 6, the model reaches near-optimal performance by the 10th epoch. The trained model demonstrated robust accuracy across three sampling intervals (PTR: $u=1$ minute, PTR: $u=2$ minutes, and PTR: $u=4$ minutes), confirming that PTR can generalize effectively to sparse trajectories with varying sampling intervals. Furthermore, after fine-tuning at each sampling interval, the performance shows minimal change in performance by the 3rd epoch, further proving its fast convergence speed.
@@ D.2. Hyperparameter Study
D.2.1. The Number of Reference Tokens $K$
We set the parameter $K$ from the set $\{128,256,512,1024\}$ to explore its impact on trajectory recovery. The experimental results are shown in Figure 7. As $K$ increases, the effectiveness of trajectory recovery also improves. This suggests that having a larger token space aids in accurately representing trajectory features. When $K$ is larger than 512, the performance improvement of PTR becomes marginal, while the computational cost will increase. To balance the performance and efficiency of the model, we set $K$ to 512.
D.2.2. The Rank $r$ of LoRA
@@
D.2.3. The Loss Weight $\lambda$
@@
D.3. Case Study
We conduct a case study on the Chengdu dataset to visualize the trajectory recovery performance with various baselines. As shown in Figure 10, we draw the truth and recovered trajectory, where red points represent the truth trajectory points and blue points indicate the recovered trajectory points. We find that the recovered trajectory aligns well with the road network, demonstrating the effectiveness of using road segment and moving rate to represent trajectory point. To facilitate a more intuitive comparison of recovery performance among different models, we focus on two distinct regions, labeled as A and B, for visual analysis. In region A, characterized by a relatively simple road network structure, all models can accurately recover road segments. Among them, PTR maps trajectory points better by leveraging PLMs. In contrast, region B exhibits a more intricate road network with multiple accessible routes. RNTrajRec captures spatial-temporal correlations of trajectories that are not adequate and recovers incorrect road segments. PTR not only excels in road segment recovery but also in accurately matching the actual trajectory points. This success can be attributed to its ability to model missing trajectory points through the implicit trajectory prompt, thereby introducing more valuable information and improving performance.
D.4. Computational Cost
We use an NVIDIA RTX A4000 card to conduct cost analysis on the Chengdu and Porto datasets with a 2-minute sampling interval by considering model size, training time, and inference time. As shown in Table 6, compared to the end-to-end deep learning model, PTR has the largest number of parameters due to its utilization of a pre-trained language model based on the BERT-small framework. Despite having a greater number of parameters, PTR exhibits quicker training and testing times as it can output results for all trajectory points at once, and the majority of parameters are frozen, unlike other models that necessitate autoregressive generation. It significantly accelerates both model training and inference processes.
@@ D.4. Computational Cost
PTR 176 / 175 24.18 / 54.14 2.56 / 10.38

### v2

@@ 5.8. Ablation Study
PLMTrajRec - Randomly initialized BERT 74.11 78.62 80.83 288.4 831.9
PLMTrajRec 81.76 84.31 88.38 187.4 366.2
As shown in Table 5 and 8, each component of PLMTrajRec is essential. The IF-guided explicit trajectory prompt provides task and trajectory-related textual information, enabling PLMs to effectively understand trajectory features. Omitting it leads to a performance decrease. When the AF-guided implicit trajectory prompt is removed, many missing locations are replaced by the ‘[MASK]’ token in BERT, limiting the model’s ability to capture road conditions for missing points, which results in suboptimal performance. For the trajectory point embedder, excluding road network for observed points causes the model to rely solely on latitude and longitude, neglecting the relationship between trajectories and the road network, which demonstrates that road network is crucial for modeling trajectory information. Additionally, removing reference tokens results in a performance decline, as PLMs are unable to directly interpret trajectory data. Using the randomly initialized BERT in PLMTrajRec has poor performance, indicating the efficiency of the PLM for trajectory recovery.
@@ 6. Conclusion
In this paper, we propose PLMTrajRec, a novel model for trajectory recovery leveraging PLM. PLMTrajRec effectively recovers trajectory even with limited dense trajectory data available, demonstrating strong scalability. The model equippes with an interval and feature-guided explicit trajectory prompt and interval-aware trajectory prompt enabling it to effectively generalize to different sampling intervals. Additionally, we introduce an area flow-guided implicit trajectory prompt to gather traffic flows in each region, and propose a road condition passing mechanism to infer missing-point conditions from nearby observations. Experimental results on two datasets with three different sampling intervals validate the effectiveness, scalability, and generalizability of the proposed model.
Appendix A Baseline Setting
We choose the following 12 methods as baselines, including five free space trajectory recovery and seven map-matched trajectory recovery.
A.1. Free-space Trajectory Recovery
@@
A.2. Map-matched Trajectory Recovery
@@
  • T2vec (Li et al., 2018) is a deep learning model for trajectory similarity learning. We use its encoder to embed the sparse trajectory.
  • T3s (Yang et al., 2021) uses LSTM and attention mechanisms to encode sparse trajectory data effectively.
  • TERI (Chen et al., 2023a) assumes that the number of missing trajectory points is unknown and proposes a two-stage trajectory recovery framework. In the first stage, a transformer-based model predicts the number of points to be recovered, and in the second stage, the same framework is used for recovery trajectory coordinates. Here, we utilize only the second stage of TERI.
  • TrajBERT (Si et al., 2023) employs a transformer encoder and a forward and backward neighbor selector to learn complex mobility patterns bi-directionally from sparse trajectories.
@@
Notably, despite T2vec, T3S, TERI, and TrajBERT employing different techniques to capture sparse trajectories’ spatial-temporal dependencies, they cannot directly generate the desired format for missing points. Therefore, after these models obtain the trajectory embeddings, we append the Decoder part of MTrajRec to output the road segment and moving ratio of the trajectory point. Denoted by T2v + Decoder, T3S + Decoder, TERI + Decoder, and TrajBERT + Decoder, respectively.
Appendix B Evaluation Metrics
@@
Appendix C Example of the IF-guided Explicit Trajectory Prompt
Consider a sparse trajectory $\mathcal{T}=\langle p_{1},\cdots,p_{N}\rangle$ of $N$ trajectory points with a sampling interval of 4 minutes, starting at 8 o’clock on Saturday and ending at 9 o’clock on Saturday. Our goal is to recover it within a sampling interval of 15 seconds. Therefore, the trajectory prompts that are related to sampling intervals are: Task Part: Sparse trajectory recovery. Target Part: Output the road segment and moving ratio for each point in the trajectory. Content Part: The sparse trajectory is sampled every four minutes and aims to recover trajectory every fifteen seconds. The movement feature-related trajectory prompts are: Time Part: The trajectory started at eight o’clock on Saturday and ended at nine o’clock on Saturday. Movement Part: Total time cost: sixty minutes zero seconds. Total space transfer distance: $z$ kilometers. Here $z={\textstyle\sum_{i=2}^{N}}\text{dist}(p_{i},p_{i-1})$, where $\mathrm{dist}(\cdot,\cdot)$ is used to calculate the distance between two trajectory points.
Appendix D Insight about the Learnable Fourier Features
Consider two trajectory points $x$ and $y$, and the feature mapping function $\Phi(x)=W_{\Phi}[\mathrm{cos}\text{ }xW_{r}||\mathrm{sin}\text{ }xW_{r}]$ in Learnable Fourier Features. The relative information $x-y$ between points $x$ and $y$ can be captured through multiplication operations, i.e.:
(19) $\begin{split}\Phi(x)\cdot\Phi(y)&=W_{\Phi}[\mathrm{cos}\text{ }xW_{r}||\mathrm{sin}\text{ }xW_{r}]\cdot W_{\Phi}[\mathrm{cos}\text{ }yW_{r}||\mathrm{sin}\text{ }yW_{r}]\\ &=||W_{\phi}||_{2}\cdot(\mathrm{cos}\text{ }x\cdot\mathrm{cos}\text{ }y+\mathrm{sin}\text{ }x\cdot\mathrm{sin}\text{ }y)\cdot||W_{r}||_{2}\\ &=||W_{\phi}||_{2}\cdot\mathrm{cos}(x-y)\cdot||W_{r}||_{2}\end{split}$
In our PLMTrajRec, after feature conversion, we input the trajectory feature into a pre-language training model based on BERT. Since there are a lot of multiplication-based attention operations in the PLM, the relative information $x-y$ can be easily modeled and utilized. This relative information helps infer crucial details like the distance between trajectory points, which is important for understanding movement. For instance, a larger distance between two points may indicate a higher likelihood of vehicle acceleration.
Appendix E The Detailed Description of Variant
  • w/o IF-guided explicit trajectory prompt: We remove the IF-guided explicit trajectory prompt.
  • w/o AF-guided implicit trajectory prompt: We use the ‘[MASK]’ token in the BERT to represent the missing location.
  • w/o dual trajectory prompts: We remove both the IF-guided explicit and implicit trajectory prompt, and the missing points are represented by the ‘[MASK]’ token in BERT.
  • PLMTrajRec - Randomly initialized BERT: We randomly initialize the parameter of BERT instead of using the pre-trained BERT based on large-scale corpus datasets.
Appendix F Appendix Experiments
@@ F.1. Hyperparameter Study
Table 7. Scalability analysis on Chengdu dataset with 1 minute and 4 minutes sampling intervals.
Sampling Interval Data Ratio 20% 40% 60% 80% 100%
[…]
RNTrajRec 60.80 998.0 62.85 931.3 64.19 909.1 65.46 835.0 67.66 886.0
MM-STGED 64.61 935.0 67.76 881.8 68.53 853.7 69.95 825.3 70.64 829.7
PLMTrajRec 68.30 585.3 71.57 515.6 72.74 512.5 73.56 488.3 74.12 483.0
RNTrajRec 75.65 846.2 79.48 784.8 79.61 769.3 81.74 742.5 81.88 702.5
MM-STGED 76.02 857.7 79.86 734.7 82.25 676.4 83.44 663.0 84.26 633.5
PLMTrajRec 81.62 397.5 84.59 352.0 85.75 328.7 86.73 305.4 87.17 290.8
F.1.1. The Number of Reference Tokens $K$
We set the parameter $K$ from the set $\{128,256,512,1024\}$ to explore its impact on trajectory recovery. The experimental results are shown in Figure 5. As $K$ increases, the effectiveness of trajectory recovery also improves. This suggests that having a larger token space aids in accurately representing trajectory features. When $K$ is larger than 512, the performance improvement of PLMTrajRec becomes marginal, while the computational cost will increase. To balance the performance and efficiency of the model, we set $K$ to 512.
F.1.2. The Rank $r$ of LoRA
@@
F.1.3. The Loss Weight $\lambda$
@@
F.2. Scalability Study on 1-minute and 4-minute Sampling Interval
The result of the scalability study on the sparse trajectory with 1-minute and 4 minutes, as shown in Table 7, the conclusion is consistent with the 2-minute sampling interval. The experiment results prove the scalability of PLMTrajRec.
Table 8. Ablation study on Porto dataset with sampling intervals at 2 minutes.
[…]
w/o IF-guided explicit trajectory prompt 66.13 66.19 81.98 157.9 314.6
w/o AF-guided implicit trajectory prompt 66.18 66.22 82.01 149.6 298.1
w/o dual trajectory prompts 65.52 65.89 81.51 171.0 339.4
w/o road network 66.11 66.27 81.90 150.7 302.8
w/o reference tokens 65.89 65.07 80.57 165.4 327.7
PLMTrajRec - Randomly initialized BERT 62.06 64.18 79.26 179.3 413.9
PLMTrajRec 66.40 66.52 82.14 141.9 294.6
F.3. Ablation Study on the Porto
The result of the ablation study on the Porto dataset, as shown in Table 8, the conclusion is consistent with the Chengdu dataset. Each component contributes to PLMTrajRec’s performance, and removing any component results in a decline in model effectiveness.
@@ F.3. Ablation Study on the Porto
PLMTrajRec 176 / 175 24.18 / 54.14 2.56 / 10.38
F.4. Computational Cost
We use an NVIDIA RTX A4000 card to conduct cost analysis on the Chengdu and Porto datasets with a 2-minute sampling interval by considering model size, training time, and inference time. As shown in Table 9, compared to the end-to-end deep learning model, PLMTrajRec has the largest number of parameters due to its utilization of a pre-trained language model based on the BERT-small framework. Despite having a greater number of parameters, PLMTrajRec exhibits quicker training and testing times as it can output results for all trajectory points at once, and the majority of parameters are frozen, unlike other models that necessitate autoregressive generation. It significantly accelerates both model training and inference processes.
F.5. Case Study
We conduct a case study on the Chengdu dataset to visualize the trajectory recovery performance with various baselines. As shown in Figure 8, we draw the truth and recovered trajectory, where red points represent the truth trajectory points and blue points indicate the recovered trajectory points. We find that the recovered trajectory aligns well with the road network, demonstrating the effectiveness of using road segment and moving rate to represent trajectory point. To facilitate a more intuitive comparison of recovery performance among different models, we focus on two distinct regions, labeled as A and B, for visual analysis. In region A, characterized by a relatively simple road network structure, all models can accurately recover road segments. Among them, PLMTrajRec maps trajectory points better by leveraging PLMs. In contrast, region B exhibits a more intricate road network with multiple accessible routes. RNTrajRec captures spatial-temporal correlations of trajectories that are not adequate and recovers incorrect road segments. PLMTrajRec not only excels in road segment recovery but also in accurately matching the actual trajectory points. This success can be attributed to its ability to model missing trajectory points through the implicit trajectory prompt, thereby introducing more valuable information and improving performance.